OpenStudy (chycora):

3. (1 pt) Janet pays $350 annually for her gym membership plus $8 each time she takes a pilates class. Last year, Janet took 20 pilates classes. What percent of her total gym cost was represented by her pilates classes? Round the answer to the nearest tenth of a percent. A. 2.3% B. 5.7% C. 31.4% D. 45.7%

OpenStudy (igreen):

First find out the cost of her pilates classes. 8 * 20 = ?

OpenStudy (chycora):

160

OpenStudy (igreen):

Yes, now divide 160 / 350

OpenStudy (igreen):

And multiply by 100

OpenStudy (chycora):

45.7 right

OpenStudy (igreen):

Yep, you got it.

OpenStudy (chycora):

thanx

OpenStudy (chycora):

@iGreen it was wrong the answer is 31.4%

OpenStudy (igreen):

Ohhhh..sorry. We were supposed to divide it to the total.. The total will be 160 + 350 = 510 160 / 510 = 0.314 = 31.4%

OpenStudy (igreen):

Sorry \(\Huge\ddot\frown\)

OpenStudy (chycora):

ur ok ur the best anyway thanx still -love